OpenStudy (anonymous):

Solve for x. 3(3x - 1) + 2(3 - x) = 0

OpenStudy (luigi0210):

3(3x-1)+2(3-x)=0 Distribute: 9x-3+6-2x Combine like terms: 7x+3

OpenStudy (luigi0210):

7x-3=0 Add 3 7x=3 Divide by 7 and you get..?

OpenStudy (anonymous):

divide what by 7?

OpenStudy (luigi0210):

Oos, I meant 7x+3=0 7x=-3 and divide -3 by 7